OpenStudy (mr.walrus):

Of the Comic books available at the comic book store, 42% are Marvel Comics, 23% are Special Edition Comics and 12% are Marvel and Special Edition Comics. A comic book will be selected randomly from the comic book store. What is the probability the comic will not be a Marvel Comic? What is the probability the selected comic will be a Special Edition given that it is a Marvel Comic? Question 2 options: 0.58 and 0.29 0.58 and 0.23 0.42 and 0.12 0.23 and 0.42

OpenStudy (rayaz):

\[P(M)=0.42 \ , P(S)=0.23 \ , P(M\ and \ S)=0.12\] \[P(Book \ isn't \ a \ marvel \ book)= 1 - P(M)\ \therefore\ 1-0.42=0.58\] \[P(S|M)=\frac{ P(M \ and \ S) }{ P(M) } \therefore\ \frac{ 0.12 }{ 0.42 }=\frac{ 2 }{ 7 }\approx0.29\]

OpenStudy (mr.walrus):

ThankYou so much!

OpenStudy (rayaz):

you're welcome